# Two-Edge Connectivity via Pac-Man Gluing

### v1

@@ Abstract
Our improvement is based on two main innovations: First, we reduce solving the problem on general graphs to solving it on structured graphs with high vertex connectivity. This high vertex connectivity ensures the existence of a 4-matching across any bipartition of the vertex set with at least 10 vertices in each part. Second, we exploit this property in a later gluing step, where isolated 2-edge-connected components need to be merged without adding too many edges. Using the 4-matching property, we can repeatedly glue a huge component (containing at least 10 vertices) to other components. This step is reminiscent of the Pac-Man game, where a Pac-Man (a huge component) consumes all the dots (other components) as it moves through a maze. These two innovations lead to a significantly simpler algorithm and analysis for the gluing step compared to the previous best approximation algorithm, which required a long and tedious case analysis.
@@ 1 Introduction
We study the 2-edge-connected spanning subgraph (2-ECSS) problem, which is a fundamental fault-tolerant network design problem. In the 2-ECSS problem, we are given a graph $G$, and the objective is to compute a connected subgraph $H$ of $G$ with the minimum number of edges such that $H$ is spanning (i.e., $V(H)=V(G)$) and $H$ is 2-edge-connected (i.e., $H$ remains connected upon the deletion of any single edge: for all edges $e\in E(H)$, $H\setminus\{e\}$ is connected), if such an $H$ exists. While computing a (1-edge) connected spanning subgraph with the minimum number of edges—a spanning tree—admits simple polynomial-time algorithms, the 2-ECSS problem is MAX-SNP hard [13, 15]. In particular, assuming $\mathrm{P}\neq\mathrm{NP}$, there is no PTAS for the 2-ECSS problem. Consequently, there has been a quest to obtain good polynomial-time approximations for the problem.
A $2$-approximation for the $2$-ECSS problem is easily achieved by augmenting a DFS tree by picking the highest back edges for each non-root vertex. The first non-trivial approximation achieving a ratio of $\nicefrac{{3}}{{2}}$ was provided by Khuller and Vishkin [26]. Subsequently, the approximation ratio was first improved to $\nicefrac{{17}}{{12}}$ by Cheriyan, Sebö, and Szigeti [10], and later to $\nicefrac{{4}}{{3}}$ independently by Hunkenschröder, Vempala, and Vetta [37, 24] and Sebö and Vygen [34] using very different techniques. Several failed attempts (see Section F of [18]) were made to improve the approximation ratio of $\nicefrac{{4}}{{3}}$, making it a natural approximation barrier. In a breakthrough, Garg, Grandoni, and Jabal Ameli [18] obtained an approximation ratio of $\nicefrac{{118}}{{89}}+\varepsilon<1.326$. Subsequently, Kobayashi and Noguchi [27] observed that if a minimum $2$-edge cover that does not contain any isolated triangles can be computed in polynomial-time, then in fact the analysis of [18] will result in an approximation ratio of $1.3+\varepsilon$. They were able to show that such a triangle-free $2$-edge cover is polynomial-time computable using a polynomial-time computable maximum triangle-free 2-matching, a result proved in Hartvigsen’s PhD thesis [22], which was recently published in a journal [23] and subsequently improved and simplified by Paluch [32]. Additionally, a PTAS for triangle-free $2$-matching is available [4], which also results in a $(1.3+\varepsilon)$-approximation for the 2-ECSS problem.
@@ 1.1 Our Result
For all $\varepsilon>0$, there exists a deterministic polynomial-time approximation algorithm for the $2$-ECSS problem with an approximation ratio of $\frac{5}{4}+\varepsilon$.
@@
The previous best approximation ratio of $1.3+\varepsilon$ [27] depends on [18], a long and a highly involved paper consisting of tedious case analysis, making the result difficult to verify. In this work, we address this shortcoming by providing a simpler proof of a better approximation ratio of $\nicefrac{{5}}{{4}}+\varepsilon$.
1.2 Our Technique
@@
Our proof is broadly based on the well-established template for such problems (as used in some form or another in [6, 7, 37, 24, 14, 29, 3]). It consists of four main steps: reducing the input instance to a structured instance, computing in polynomial time an object $\mathcal{O}$ (a minimum 2-edge cover with no isolated triangles) of the structured instance that is closely related to and costs less than an optimal desired object (minimum 2-ECSS), then converting $\mathcal{O}$ into a feasible solution (2-ECSS) by first covering the bridges in each connected component, followed by gluing the various connected components together while keeping the additional cost of the conversion to a small factor $c$ of the cost of $\mathcal{O}$. This results in a $(1+c)$-approximation. We will elaborate on these steps in detail in the following sections.
@@
In brief, for the first three steps, we closely follow [18], which is an uncomplicated part of their paper, along with [27] for the second step, which helps in eliminating isolated triangles, an important fact without which the gluing would be cumbersome. Our main insight in this part is to make the input even more structured than in [18] by getting rid of all large 3-vertex cuts. This allows us to obtain a 4-matching lemma, which states that for any vertex partition of a structured input into two parts consisting of at least 10 vertices each, there exists a matching of size 4 going between the two parts. This lemma is then exploited in the gluing step. In the gluing step, we first glue a few components to obtain a huge component (consisting of at least 10 vertices). Then, with the aid of the 4-matching lemma, we are repeatedly able to glue a huge component with the other components, while keeping $c$ to be $\nicefrac{{1}}{{4}}$. This step is reminiscent of the popular Pac-Man game, where a Pac-Man (a huge component) eats up all the dots (other components) as it moves through a maze (component graph). Like in [18], we lose an additional $\varepsilon$ approximation factor in the reduction step itself, giving us an approximation ratio of $1+\nicefrac{{1}}{{4}}+\varepsilon=\nicefrac{{5}}{{4}}+\varepsilon$.
@@ 1.3 Additional Related Work
The 2-ECSS problem is closely related to augmentation problems, where we are given a graph having a specific structure (e.g., spanning tree, matching, forest, cactus, or disjoint paths) along with a set of additional edges called links. The task is to add the minimum number of links to the structure so that the resulting graph becomes spanning and 2-edge connected. For the (spanning) tree augmentation problem, several better than $2$-approximations are known [1, 5, 8, 9, 12, 14, 16, 21, 28, 29, 30, 31, 35, 36], with 1.393 as the currently best known approximation ratio by Cecchetto et al. [5]. For the forest augmentation problem, the first non-trivial approximation was obtained by Grandoni et al. [20] with a ratio of 1.9973. For the matching augmentation problem the current best approximation ratio is $\nicefrac{{13}}{{8}}$ [19] by Garg et al., while other non-trivial approximations are also known [2, 6, 7]. All the above problems have a natural weighted version. For the weighted 2-ECSS problem, obtaining an approximation ratio better than 2 is a major open problem. For the weighted tree augmentation problem the current best approximation ratio is $1.5+\varepsilon$ by Zenklusen and Traub [36]. The k-ECSS problem for $k\geq 2$ is another natural generalization of the 2-ECSS problem (see for e.g., [11, 17]).
@@ 1.4 Organization of the Paper
The paper proceeds as follows. In Section 2 we start with some preliminaries consisting of some basic definitions and notations. Subsequently, we tackle the four main steps of our approach in the following four sections: reduction to structured graphs in Section 3, obtaining a canonical 2-edge cover in Section 4, bridge covering in Section 5, and finally, we present the gluing step, the crux of this work, in Section 6. Certain proofs have been moved to the Appendix.
@@ 2 Preliminaries
We use standard graph theory notation. For a graph $G$, $V(G)$ and $E(G)$ refers to its vertex and edge sets, respectively. We use $\deg_{G}(v)$ to represent the degree of a vertex $v$ in the graph $G$ (when $G$ is clear from the context, we sometimes omit the subscript $G$). A pendant vertex in a graph is a vertex that is adjacent to exactly one other vertex. For a graph $G$, and a vertex set $S\subseteq V(G)$, $G[S]$ refers to the subgraph of $G$ induced on the vertex set $S$. We use components of a graph to refer to its maximal connected components. A cut vertex in a graph refers to a vertex in the graph, so that deleting it results in increasing the number of components of the graph. A $k$-vertex cut in a graph for $k\geq 1$ refers to a set of $k$ vertices such that deleting them results in increasing the number of components of the graph. A graph with no cut vertex is called 2-vertex connected (2VC). Similarly, a bridge in a graph refers to an edge in the graph, so that deleting it results in increasing the number of components in the graph. A connected graph without bridges is called 2-edge connected (2EC). We say a subgraph $H$ of a graph $G$ is spanning if $V(H)=V(G)$.
A spanning subgraph $H$ of a graph $G$ is called a 2-edge cover of $G$ if the degree of each vertex in $H$ is at least 2. A cycle refers to a simple cycle. We use $\mathcal{C}_{i}$ to represent a cycle on $i$ vertices and refer to $\mathcal{C}_{3}$ as a triangle. We say a 2-edge cover is triangle-free, if none of its components is a triangle.
@@ 3 Reduction to Structured Graphs
In our approximation algorithm for the 2-ECSS problem we will assume that the input is 2EC, otherwise it has no feasible solutions, which is easy to check: delete each edge and check if the graph is connected. Note that the input graph may contain parallel edges or self loops. In [18], structured graphs were defined, and it was shown that, for each small enough $\varepsilon>0$, if the 2-ECSS problem can be approximated in polynomial time on structured graphs with an approximation ratio $\alpha\geq\nicefrac{{6}}{{5}}$, then it can be approximated in polynomial time on general graphs with only a $2\varepsilon$ loss in approximation, i.e., with an approximation ratio of $\alpha+2\varepsilon$. In this work, we define structured graphs to be even more restricted than in [18], and prove a similar approximation-preserving reduction up to an $O(\varepsilon)$ loss.
Informally, a structured graph is a simple graph that is 2VC with sufficiently many vertices and does not contain any contractible subgraphs, irrelevant edges, non-isolating 2-vertex cuts, or large 3-vertex cuts. Forbidding large 3-vertex cuts in the definition is the only difference compared to the definition in [18]. We now define the various terms appearing in the definition of structured graphs before defining structured graphs formally.
Definition 1 ($\alpha$-contractible subgraph [18]).
@@
The intuition behind defining $\alpha$-contractible subgraphs is the following. While designing an $\alpha$-approximation algorithm for the 2-ECSS problem, if we can find an $\alpha$-contractible subgraph $C$ in polynomial-time, we can contract $C$ into a single vertex and find an $\alpha$-approximate solution on $G|C$, and add the edges of $C$ to get an $\alpha$-approximate solution for $G$. The approximation guarantee holds since $\mathsf{opt}(G)\geq\mathsf{opt}(G|C)+\mathsf{opt}(C)$, which is true since $\mathsf{OPT}(G)$ needs to have at least $\nicefrac{{1}}{{\alpha}}\cdot\lvert E(C)\rvert$ from $G[V(C)]$. Thus, in our approximation-preserving reduction all constant-sized $\alpha$-contractible subgraphs will be identified and handled.
Definition 2 (irrelevant edge [18]).
@@
Definition 3 (non-isolating 2-vertex cut [18]).
Given a graph $G$, and a 2-vertex cut $\{u,v\}\subseteq V(G)$ of $G$, we say that $\{u,v\}$ is isolating if $G\setminus\{u,v\}$ has exactly two connected components, one of which is of size 1. Otherwise, it is non-isolating.
@@
Definition 5 (($\alpha,\varepsilon)$-structured graph).
@@
    $\alpha$-contractible subgraphs of size at most $\frac{2}{\varepsilon}$,
@@
In [18], it was shown how to get rid of parallel edges, self loops, irrelevant edges, cut-vertices, small $\alpha$-contractible subgraphs and non-isolating 2-vertex cuts via an approximation-preserving reduction. We add to their proof and show that we can also get rid of large 3-vertex cuts by generalizing their proof for non-isolating 2-vertex cuts in various ways. This is a significant part of this work. In Appendix A we prove the following approximation-preserving reduction. This proof is a bit long, but it admits a systematic case analysis. After understanding the main ideas through the first few cases, the remaining cases can actually be left as a simple exercise (though for completeness, we cover all the cases). {restatable}lemmalemmaReduction For all $\alpha\geq\frac{5}{4}$ and $\varepsilon\in(0,\frac{1}{24}]$, if there exists a deterministic polynomial-time $\alpha$-approximation algorithm for $2$-ECSS on ($\alpha,\varepsilon)$-structured graphs, then there exists a deterministic polynomial-time $(\alpha+4\varepsilon)$-approximation algorithm for $2$-ECSS.
@@
Benefits of stuctured graphs. After we have the above reduction, we need to design approximation algorithms only for structured graphs. How do we exploit the fact that the input graph is structured? We will make generous use of the fact that there are no small contractible subgraphs. Furthermore, it was shown in [18] that certain 3-matchings exist in a structured graph, which will be very useful.
Lemma 2 ($3$-Matching Lemma [18]).
Let $G=(V,E)$ be a $2$VC simple graph without irrelevant edges and without non-isolating 2-vertex cuts. Consider any partition $(V_{1},V_{2})$ of $V$ such that for each $i\in\{1,2\}$, $|V_{i}|\geq 3$, and if $|V_{i}|=3$, then $G[V_{i}]$ is a triangle. Then, there exists a matching of size $3$ between $V_{1}$ and $V_{2}$.
We further make use of the non-existence of large $3$-vertex cuts in structured graphs to establish a $4$-matching-lemma, which is going to be very useful in the gluing step. [^1] In fact, with similar techniques we can also get rid of large $k$-vertex cuts (for any constant $k$), where both sides of the partition have $\Omega(\frac{k}{1-\alpha})$ many vertices, leading to a similar $k$-Matching Lemma. However, we do not see an easy way to exploit this fact in improving the approximation ratio with the current techniques.
@@
Consider the bipartite graph $F$ induced by the edges with exactly one endpoint in $V_{1}$ (and the other in $V_{2}$). For contradiction, we assume that a maximum matching $M$ in $F$ has size $\lvert M\rvert\leq 3$. Then, by the König Egerváry Theorem (see, e.g., [33]), there exists a vertex cover $C$ of $F$ of size $\lvert M\rvert$. Since $C$ is a vertex cover of $F$, there are no edges in $F$ (hence in $G$) between $U_{1}\coloneqq V_{1}\setminus C$ and $U_{2}\coloneqq V_{2}\setminus C$. Therefore, $C$ is a 3-vertex cut in $G$, which is large as $\lvert U_{1}\rvert,\lvert U_{2}\rvert\geq 7$ (because $\lvert V_{1}\rvert,\lvert V_{2}\rvert\geq 10$ and $\lvert C\rvert\leq 3$), a contradiction. ∎

### v2

@@ Abstract
Our improvement is based on two main innovations: First, we reduce solving the problem on general graphs to solving it on structured graphs with high vertex connectivity. This high vertex connectivity ensures the existence of a 4-matching across any bipartition of the vertex set with at least 10 vertices in each part. Second, we exploit this property in a later gluing step, where isolated 2-edge connected components need to be merged without adding too many edges. Using the 4-matching property, we can repeatedly glue a huge component (containing at least 10 vertices) to other components. This step is reminiscent of the Pac-Man game, where a Pac-Man (a huge component) consumes all the dots (other components) as it moves through a maze. These two innovations lead to a significantly simpler algorithm and analysis for the gluing step compared to the previous best approximation algorithm, which required a long and tedious case analysis.
@@ 1 Introduction
We study the 2-edge-connected spanning subgraph (2-ECSS) problem, which is a fundamental fault-tolerant network design problem. In the 2-ECSS problem, we are given a graph $G$, and the objective is to compute a connected subgraph $H$ of $G$ with the minimum number of edges such that $H$ is spanning (i.e., $V(H)=V(G)$) and $H$ is 2-edge-connected (i.e., $H$ remains connected upon the deletion of any single edge: for all edges $e\in E(H)$, $H\setminus\{e\}$ is connected), if such an $H$ exists. While computing a (1-edge) connected spanning subgraph with the minimum number of edges—a spanning tree—admits simple polynomial-time algorithms, the 2-ECSS problem is MAX-SNP hard [CL99, Fer98]. In particular, assuming $\mathrm{P}\neq\mathrm{NP}$, there is no PTAS for the 2-ECSS problem. Consequently, there has been a quest to obtain good polynomial-time approximations for the problem.
A $2$-approximation for the $2$-ECSS problem is easily achieved by augmenting a DFS tree by picking the highest back edges for each non-root vertex. The first non-trivial approximation achieving a ratio of $\nicefrac{{3}}{{2}}$ was provided by Khuller and Vishkin [KV92]. Subsequently, the approximation ratio was first improved to $\nicefrac{{17}}{{12}}$ by [CSS01], and later to $\nicefrac{{4}}{{3}}$ independently by Hunkenschröder, Vempala, and Vetta [VV00, HVV19] and [SV14] using very different techniques. Several failed attempts (see Section F of [GGA23]) were made to improve the approximation ratio of $\nicefrac{{4}}{{3}}$, making it a natural approximation barrier. In a breakthrough, Garg, Grandoni, and Jabal Ameli [GGA23] obtained an approximation ratio of $\nicefrac{{118}}{{89}}+\varepsilon<1.326$. Subsequently, Kobayashi and Noguchi [KN23] observed that if a minimum $2$-edge cover that does not contain any isolated triangles can be computed in polynomial-time, then in fact the analysis of [GGA23] will result in an approximation ratio of $1.3+\varepsilon$. They were able to show that such a triangle-free $2$-edge cover is polynomial-time computable using a polynomial-time computable maximum triangle-free 2-matching, a result proved in Hartvigsen’s PhD thesis [Har84], which was recently published in a journal [Har24] and subsequently improved and simplified by Paluch [Pal23]. Additionally, PTASs for triangle-free $2$-matching are available [BGA23a, KN25] (with the proof by [KN25] being only a few pages), which also result in a $(1.3+\varepsilon)$-approximation for the 2-ECSS problem.
@@ 1.1 Our Result
For all $\varepsilon>0$, there exists a deterministic polynomial-time approximation algorithm for the $2$-ECSS problem with an approximation ratio of $\nicefrac{{5}}{{4}}+\varepsilon$.
@@
The previous best approximation ratio of $1.3+\varepsilon$ [KN23] depends on [GGA23], a long and highly involved paper consisting of a tedious case analysis, making the result difficult to verify. In this work, we address this shortcoming by providing a simpler, shorter, and modular proof of a better approximation ratio of $\nicefrac{{5}}{{4}}+\varepsilon$.
1.2 Our Techniques
@@
Our proof is broadly based on the well-established template for such problems (as used in some form or another in [Che+20, Che+23, GHM23, VV00, HVV19, Eve+09, KN18, BGA23]). It consists of four main steps: reducing the input instance to a structured instance, computing in polynomial time an object $\mathcal{O}$ (a minimum 2-edge cover with no isolated triangles) of the structured instance that is closely related to and costs less than an optimal desired object (minimum 2-ECSS), then converting $\mathcal{O}$ into a feasible solution (2-ECSS) by first covering the bridges in each connected component, followed by gluing the various connected components together while keeping the additional cost of the conversion to a small factor $c$ of the cost of $\mathcal{O}$. This results in a $(1+c)$-approximation. We will elaborate on these steps in detail in the following sections.
@@
In brief, for the first three steps, we closely follow [GGA23], which are uncomplicated parts of their paper, along with [KN23] for the second step, which helps in eliminating isolated triangles, an important fact without which the gluing would be cumbersome. Our main insight in this part is to make the input even more structured than in [GGA23] by getting rid of all large 3-vertex cuts. This allows us to obtain a 4-matching lemma, which states that for any vertex partition of a structured input into two parts consisting of at least 10 vertices each, there exists a matching of size 4 going between the two parts. This lemma is then exploited in the gluing step.
In the gluing step, we first glue a few components to obtain a huge component (consisting of at least 10 vertices). Then, with the aid of the 4-matching lemma, we are repeatedly able to glue a huge component with the other components, while keeping $c$ to be $\nicefrac{{1}}{{4}}$. This step is reminiscent of the popular Pac-Man game, where a Pac-Man (a huge component) eats up all the dots (other components) as it moves through a maze (component graph). An important advantage of our approach is that the gluing step, which has traditionally been considered the most complex and intricate part of the process (as seen in [GGA23]), is now notably simplified, making it as straightforward as the other steps in the proof. Like in [GGA23], we lose an additional $\varepsilon$ approximation factor in the reduction step itself, giving us an approximation ratio of $1+\nicefrac{{1}}{{4}}+\varepsilon=\nicefrac{{5}}{{4}}+\varepsilon$.
Most of the proofs for the first three steps have been revisited from [GGA23] or are slight extensions, and they have been moved to the appendix for completeness. The primary new contribution in the appendix is the preprocessing step involving the elimination of large 3-vertex cuts.
In summary, our technique separates the technical complications of the gluing used in [GGA23] into two modular and independent components, the reduction to graphs without large 3-vertex cuts and the Pac-Man gluing. We believe that the simplicity and modularity of our approach makes it applicable in and relevant for future work. This streamlined approach transforms the gluing step, once the most challenging part, into an intuitive and manageable procedure.
@@ 1.3 Additional Related Work
Survivable Network Design.
The 2-ECSS problem is closely related to augmentation problems, where we are given a graph having a specific structure (e.g., spanning tree, matching, forest, cactus, or disjoint paths) along with a set of additional edges called links. The task is to add the minimum number of links to the structure so that the resulting graph becomes spanning and 2-edge connected. For the (spanning) tree augmentation problem, several better than $2$-approximations are known [Adj19, CTZ21, CG18, CG18a, CN13, Eve+09, Fio+18, GKZ18, KN16, KN18, Nag03, Nut21, TZ25], with 1.393 as the currently best known approximation ratio by [CTZ21]. For the forest augmentation problem, the first non-trivial approximation was obtained by [GAT22] with a ratio of 1.9973. For the matching augmentation problem the current best approximation ratio is $\nicefrac{{13}}{{8}}$ [GHM23] by [GHM23], while other non-trivial approximations are also known [BDS22, Che+20, Che+23]. All the above problems have a natural weighted version, and a general result of Jain [J01] provides a $2$-approximation for them. For the weighted 2-ECSS problem, obtaining an approximation ratio better than 2 is a major open problem. For the weighted tree augmentation problem the current best approximation ratio is $1.5+\varepsilon$ by Traub and Zenklusen [TZ25]. The k-ECSS problem for $k\geq 2$ is another natural generalization of the 2-ECSS problem (see for e.g., [CT00, GG12]).
Graphic TSP.
Another closely related problem is the Traveling Salesperson Problem (TSP). In metric TSP, one is given an undirected complete graph with weights on the edges that satisfy the triangle-inequality, and the goal is to find a tour through all vertices of minimum total weight. In a recent breakthrough [KKG21], Karlin, Klein, and Oveis Gharan improved the long-standing best approximation factor of $1.5$ by Christofides [Chr22] to a (1.5c)(1.5-c)-approximation, where $c>10^{-36}$. The algorithm makes use of the Help-Karp LP-relaxation of the problem, which has an integrality gap of $\nicefrac{{4}}{{3}}$, indicating a potential answer for the best-possible approximation guarantee of metric TSP. The example of this integrality gap is actually an instance of graphic TSP, which is a special case of metric TSP. Therein, we are given an undirected unweighted graph $G$ and one has to find an Eulerian multigraph while minimizing the number of edges. Sebö and Vygen [SV14] proved the currently best approximation guarantee of $1.4$ for graphic TSP. They use very similar techniques in the $1.4$-approximation for graphic TSP and the $\nicefrac{{4}}{{3}}$-approximation for $2$-ECSS, which highlights the close connection between the two problems. Therefore, our new techniques for 2-ECSS could potentially contribute to further advancements in graphic TSP.
Concurrent Work.
Shortly after the first draft of this work was made publicly available, Bosch-Calvo, Grandoni, and Jabal Ameli [BGA24] presented a different $\nicefrac{{5}}{{4}}$-approximation. While both works build on the framework of [GGA23], the approaches diverge substantially in their key innovations. In particular, [BGA24] removes the $\varepsilon$ additive term in the preprocessing phase by exploiting an observation used in earlier works [Che+20, Che+23, GHM23].
In contrast, in our approach the $\varepsilon$ term is not eliminated. Instead, it is deliberately used to remove large 3-vertex cuts in the preprocessing phase, which is crucial for establishing the 4-matching lemma. Combined with our Pac-Man-style gluing procedure, this leads to a proof that is modular, conceptually simple, and straightforward to verify—rather than relying on the long, extensive, and intricate case-analysis-based gluing employed in [BGA24].
Subsequently, the two lines of work were merged in a joint paper [Bos+25], obtaining a $\nicefrac{{5}}{{4}}$-approximation where our Pac-Man gluing technique is combined with the preprocessing of [BGA24], which significantly reduces the complexity of the gluing step. However, the 4-matching lemma from this work was not leveraged in the joint paper; incorporating it would further streamline the gluing as compared to [Bos+25], though it introduces an additional $\varepsilon$ additive term in the approximation ratio.
@@ 1.4 Organization of the Paper
The paper proceeds as follows. In Section˜2 we start with some preliminaries consisting of some basic definitions and notations. Subsequently, we tackle the four main steps of our approach in the following four sections: reduction to structured graphs in Section˜3, obtaining a canonical 2-edge cover in Section˜4, bridge covering in Section˜5, and the gluing step, which forms the core of this work, in Section˜6. For completeness, the appendix contains detailed proofs of the first three steps, most of which have been revisited from [GGA23] or are mild generalizations. The main new contribution in the appendix is the elimination of large 3-vertex cuts in the preprocessing. Finally, we conclude in Section˜7, summarizing our contributions and discussing potential avenues for future research.
@@ 2 Preliminaries
We use standard graph theory notation. For a graph $G$, $V(G)$ and $E(G)$ refers to its vertex and edge sets, respectively. We use $\deg_{G}(v)$ to represent the degree of a vertex $v$ in the graph $G$ (when $G$ is clear from the context, we omit the subscript $G$). A pendant vertex in a graph is a vertex that is adjacent to exactly one other vertex. For a graph $G$, and a vertex set $S\subseteq V(G)$, $G[S]$ refers to the subgraph of $G$ induced on the vertex set $S$. We use components of a graph to refer to its maximal connected components. A cut vertex in a graph refers to a vertex in the graph whose deletion results in increasing the number of components of the graph. A $k$-vertex cut in a graph for $k\geq 1$ refers to a set of $k$ vertices such that deleting them results in increasing the number of components of the graph. A graph that does not contain any cut vertex is called 2-vertex connected (2VC). Similarly, a bridge in a graph refers to an edge in the graph whose deletion results in increasing the number of components in the graph. A connected graph without bridges is called 2-edge connected (in short, 2EC). We say a subgraph $H$ of a graph $G$ is spanning if $V(H)=V(G)$.
A spanning subgraph $H$ of a graph $G$ is called a 2-edge cover of $G$ if the degree of each vertex in $H$ is at least 2. A cycle refers to a simple cycle. We use $\mathcal{C}_{i}$ to represent a cycle on $i$ vertices, and $\mathcal{C}_{3}$ to refer to a triangle. We say a 2-edge cover is triangle-free if none of its components is a triangle.
@@ 3 Reduction to Structured Graphs
In our approximation algorithm for the 2-ECSS problem, we will assume that the input is 2EC. Otherwise, it has no feasible solutions, which is easy to check: delete each edge and check if the graph is connected. Note that the input graph may contain parallel edges or self loops. In [GGA23], structured graphs were defined, and it was shown that, for each small enough $\varepsilon>0$, if the 2-ECSS problem can be approximated in polynomial time on structured graphs with an approximation ratio $\alpha\geq\nicefrac{{6}}{{5}}$, then it can be approximated in polynomial time on general graphs with only a $O(\varepsilon)$ loss in approximation, i.e., with an approximation ratio of $\alpha+O(\varepsilon)$. In this work, we define structured graphs to be even more restricted than in [GGA23], and yet prove a similar approximation-preserving reduction that maintains an approximation ratio of $\alpha+O(\varepsilon)$.
Informally, a structured graph is a simple graph that is 2VC with sufficiently many vertices and does not contain any contractible subgraphs, irrelevant edges, non-isolating 2-vertex cuts, or large 3-vertex cuts. Forbidding large 3-vertex cuts in the definition is the only difference compared to the definition in [GGA23]. We now define the various terms appearing in the definition of structured graphs before defining structured graphs formally.
Definition 1 ($\alpha$-contractible subgraph [GGA23]).
@@
The intuition behind defining $\alpha$-contractible subgraphs is the following: While designing an $\alpha$-approximation algorithm for the 2-ECSS problem, if we can find an $\alpha$-contractible subgraph $C$ in polynomial time, we can contract $C$ into a single vertex and find an $\alpha$-approximate solution on $G|C$, and add the edges of $C$ to get an $\alpha$-approximate solution for $G$. The approximation guarantee holds since $\alpha\cdot\mathsf{opt}\geq\alpha\cdot\mathsf{opt}(G|C)+|E(C)|$, which is true since $\mathsf{opt}(G)\geq\mathsf{opt}(G|C)+|\mathsf{OPT}[V(C)]|$ and $\mathsf{OPT}[V(C)]$ needs to contain at least $\nicefrac{{1}}{{\alpha}}\cdot\lvert E(C)\rvert$ edges of $G[V(C)]$. Thus, in our approximation-preserving reduction all constant-sized $\alpha$-contractible subgraphs will be identified and handled.
Definition 2 (irrelevant edge [GGA23]).
@@
Definition 3 (non-isolating 2-vertex cut [GGA23]).
Given a graph $G$, and a 2-vertex cut $\{u,v\}$ of $G$, we say that $\{u,v\}$ is isolating if $G\setminus\{u,v\}$ has exactly two connected components, one of which is of size 1. Otherwise, it is non-isolating.
@@
Definition 5 $($$\alpha,\varepsilon)$-structured graph).
@@
    $\alpha$-contractible subgraphs of size at most $\frac{4}{\varepsilon}$,
@@
In [GGA23], it was shown how to get rid of parallel edges, self loops, irrelevant edges, cut-vertices, small $\alpha$-contractible subgraphs, and non-isolating 2-vertex cuts via an approximation-preserving reduction. We show that we can also get rid of large 3-vertex cuts by generalizing their proof for non-isolating 2-vertex cuts in various ways. This is a significant part of this work. In Appendix˜A, we prove the following approximation-preserving reduction. This proof is a bit long, but it admits a systematic case analysis. After understanding the main ideas through the first few cases, the remaining cases can be left as a simple exercise (though for completeness, we cover all the cases, including the ones proved in [GGA23] for Properties 1-3). {restatable}lemmalemmaReduction For all $\alpha\geq\frac{5}{4}$ and $\varepsilon\in(0,\frac{1}{24}]$, if there exists a deterministic polynomial-time $\alpha$-approximation algorithm for $2$-ECSS on ($\alpha,\varepsilon)$-structured graphs, then there exists a deterministic polynomial-time $(\alpha+4\varepsilon)$-approximation algorithm for $2$-ECSS.
@@
Benefits of structured graphs. Given the above reduction, we need to design approximation algorithms only for structured graphs. How do we exploit the fact that the input graph is structured? We will make generous use of the fact that there are no small contractible subgraphs. Furthermore, it was shown in [GGA23] that certain 3-matchings exist in a structured graph, which will be very useful.
Lemma 2 ($3$-Matching Lemma [GGA23]).
Let $G=(V,E)$ be a 22VC simple graph without irrelevant edges and without non-isolating 2-vertex cuts. Consider any partition $(V_{1},V_{2})$ of $V$ such that for each $i\in\{1,2\}$, $|V_{i}|\geq 3$, and if $|V_{i}|=3$, then $G[V_{i}]$ is a triangle. Then, there exists a matching of size $3$ between $V_{1}$ and $V_{2}$.
We further make use of the non-existence of large $3$-vertex cuts in structured graphs to establish a $4$-matching-lemma, which is going to be very useful in the gluing step. In fact, with similar techniques, we can also exclude large $k$-vertex cuts (for any constant $k$), where both sides of the partition have $\Omega(\frac{k}{1-\alpha})$ many vertices, leading to a similar $k$-Matching Lemma. However, we do not see an easy way to exploit this fact in improving the approximation ratio with the current techniques.
@@
Consider the bipartite graph $F$ induced by the edges with exactly one endpoint in $V_{1}$ (and the other in $V_{2}$). For contradiction, we assume that a maximum matching $M$ in $F$ has size $\lvert M\rvert\leq 3$. Then, by the König-Egerváry Theorem (see, e.g., [Sch+03]), there exists a vertex cover $C$ of $F$ of size $\lvert M\rvert$. Since $C$ is a vertex cover of $F$, there are no edges in $F$ (hence in $G$) between $U_{1}\coloneqq V_{1}\setminus C$ and $U_{2}\coloneqq V_{2}\setminus C$. Therefore, $C$ is a 3-vertex cut in $G$, which is large because $\lvert V_{1}\rvert,\lvert V_{2}\rvert\geq 10$ and $\lvert C\rvert\leq 3$ imply $\lvert U_{1}\rvert,\lvert U_{2}\rvert\geq 7$; this is a contradiction. ∎